A person travels by car from one city to another with different constant speeds between pairs of cities. She drives for 50.0 min at 95.0 km/h, 7.0 min at 80 km/h, and 50.0 min at 55.0 km/h, and spends 30.0 min eating lunch and buying gas. '

a. determine the average speed for the trip
b.Determine the distance between the initial and final cities along this route.
1
Expert's answer
2012-09-18T11:11:13-0400
a. determine the average speed for the trip

The total distance traveled is

L = 95/60[km/m]*50[m] + 80/60[km/m]*7[m] + 55/60[km/m]*50[m] = 134+1/3 [km].

The total time elapsed is

T = 50[m] + 7[m] + +50[m] + 30[m] = 137 [m].

So, the average speed for the trip was

V = L/T = (134+1/3)[km] / (137)[m] ≈ 0.9805 [km/m] = 58.83 [km/h].
&
b.Determine the distance between the initial and final cities along this route.

The distance between the initial and final cities is L = 134+1/3 [km].